Chapter 08, Section 8.2, Problem 011 For a data set obtained from a sample, n a. What is the point estimate of H? 81 and x = 45.55. It is known that ? = 4.2. The point estimate is b. Make a 95% confidence interval for Round your answers to two decimal places. c. What is the margin of error of estimate for part b? Round your answer to three decimal places. Click if you would like to Show Work for this question: Open Show Work
Answer #1

0.11 a. \mu=\overline{x}=45.55

b. z value is 1.96, So Margin of Error=E=z*\frac{\sigma }{\sqrt{n}}=1.96*\frac{4.2}{9}=0.91

So CI=\mu \pm E=45.55 \pm 0.91=(44.64,46.46)

c. E=z*\frac{\sigma }{\sqrt{n}}=1.96*\frac{4.2}{9}=0.91

0.17a. n=(\frac{z*\sigma}{E})^2=101.89=102